# A Generative Marker Enhanced End-to-End Framework for Argument Mining

### v2

@@ III-C3 Marker Knowledge Transfer through Encoding (E-MKT)
III-C4 Discourse Marker Knowledge Transfer (D-MKT)
@@ III-D The ME-argTANL Framework
IV Experimental Setup
IV-A Datasets
We consider three standard datasets to evaluate the proposed methods. A brief description of the three datasets is given below.
[…]
@@ IV-A Datasets
IV-B Training Details
We conducted all experiments using the T5-Base model [25] on Nvidia A100 GPUs, with each reported result averaged over 5 runs [^1] The implementation will be made available upon acceptance.. The hyperparameter settings for the argTANL framework are as follows: For the CDCP dataset, a batch size of 4 is used, while for AAE and AAE-FG, the batch size is set to 8. Across all datasets, we utilize the AdamW optimizer with a learning rate of 0.0005. The maximum input length is set to 1024 tokens for CDCP and 512 tokens for AAE and AAE-FG. In the end-to-end and component-only training scenarios, we run for more than 10000 steps, with checkpoints saved every 200 steps. For the relation-only scenario, we run for 5000 steps, with checkpoints also saved every 200 steps. During inference, beam search is used with a beam length of 8 across all datasets. For all strategies (A-MKT, SM-MKT, D-MKT, and E-MKT) of Marker-Based Fine-Tuning, we use a batch size of 16, the AdamW optimizer with a learning rate of 0.0005, a maximum input length of 512 tokens, and 10 epochs of training. We use the following libraries: (i) TANL framework [^2] https://github.com/amazon-science/tanl [17], and (ii) HuggingFace’s Transformers [^3] https://github.com/huggingface [26].
IV-C Baseline Methods
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
TABLE III: Experiment results on ACRE task with comparable baselines. Best scores are marked in bold. * indicates the baseline results we produced by running their open-source code.
@@ IV-C Baseline Methods
IV-D Performance Metrics for Evaluation
After generating the target ANL, we post-process the sequence by removing the symbol tokens to get a cleaned text. Finally, for a comprehensive evaluation, AC and AR tuples are created, including their types and corresponding boundaries. Following [28], we evaluate the results with micro-F1 score for both ACE and ACRE tasks, where an exact match with the gold label is considered as a true label. But, for the ARC task in Relation-only variant, we are giving the ADU spans (without component types) as Oracle-Spans (OS) in both inputs and outputs. So, we only calculate the micro-F1 score for the generated AR labels.
V Results and Discussion
@@
Table III compares the performance of the proposed frameworks with the baselines on the ACRE task in terms of micro F1-scores for component classification and relation classification. For the AAE dataset, $T5_{ME-argTANL}$ turns out to be the best among the proposed frameworks, surpassing argTANL and all the MFS-based strategies, for component classification. While $T5_{ME-argTANL}$ outperforms several baselines for component and relation classifications, it gave a competitive performance to the current state-of-the-art models for component classification. We observe a similar performance on the AAE-FG dataset. For the AAE-FG dataset, $T5_{ME-argTANL}$ surpasses the current state-of-the-art result for component classification.
@@
TABLE IV: Performance difference on the ACE task: Component-Only vs. End-to-End variants. Among all techniques, minimum differences between these two variants are marked in bold.
@@
V-A Component Only vs End-to-End Variants
Table IV compares the performance of models on the ACE task in the two formulations. Analyzing the argumentative structure of AAE and AAE-FG datasets reveals that argument components always form a tree, indicating a single parent for each child. In this well-structured dataset, the performance of the models drops in the Component-only variant compared to the end-to-end variant. The CDCP dataset contains many isolated components. It does not necessarily form a tree but can form a graph (e.g., $t_{1}^{c}$$t_{2}^{c}$, $t_{2}^{c}$$t_{3}^{c}$, and a transitive relation $t_{1}^{c}$$t_{3}^{c}$). We observe the similar performance of models, except E-MKT, in the two variants for CDCP corpus. This result indicates that ACs and ARs benefit from mutual feature information, suggesting synergy in an end-to-end setup. For Component-only variant, the D-MKT strategy proves to be beneficial for all the datasets. When comparing the compatibility of the E-MKT with other MFS strategies, it appears more suitable for the Component-only variant rather than the end-to-end for the ACE task. It exhibits the minimum performance difference in the Component-only setup for both versions of the AAE corpus and even shows performance improvements for the CDCP corpus.
V-B Results of the Relation-Only Variant
Figure 6: Performance comparison of “with or without AC/AR types” cases for the ACRE task between AAE (coarse-grained) and AAE-FG (fine-grained) datasets.
TABLE V: Performance of proposed frameworks on Relation-only variant.
Corpus Model Micro-F1
$T5_{argTANL}$ 96.27
AAE $T5_{E-MKT}$ 96.77
or $T5_{A-MKT}$ 96.34
AAE-FG $T5_{SM-MKT}$ 96.32
$T5_{D-MKT}$ 96.47
$T5_{argTANL}$ 97.27
$T5_{E-MKT}$ 97.42
CDCP $T5_{A-MKT}$ 97.42
$T5_{SM-MKT}$ 97.35
$T5_{D-MKT}$ 97.42
Table V shows all TANL frameworks find relation classification easy if AC spans are given. All the models exhibit comparable micro-F1 scores. It highlights the capability of the T5-base to detect accurate relations between two text spans when presented with well-defined appropriate input/output templates.
V-C Coarse-Grained vs Fine-Grained Label Semantics
We take argTANL framework to perform this analysis. Table III shows a strong correlation between performance and label granularity. Most baselines perform well on the AAE dataset, where labels are coarse-grained, requiring minimal specialized understanding of context. However, all the baselines struggle with the AAE-FG and CDCP datasets. The labels in these datasets are fine-grained but at flat levels, requiring advanced model capabilities for accurate interpretation. Despite sharing identical text content, the AAE and AAE-FG datasets have different label granularity, resulting in significant performance differences. For example, in the AAE-FG or CDCP dataset, the AC labels, such as Fact or Value, correspond to Claim at a coarse-grained level. With limited context in the paragraph, the models find it relatively easy to identify the AC as a Claim compared to distinguishing it as Fact or Value. To investigate further, we analyzed both versions of the AAE dataset to determine (i) the number of ACs correctly identified with and without considering their types (referred to as ADUs) and (ii) the number of ARs correctly identified between these ADUs, with and without their AR types. Figure 6 demonstrates that when AC and AR type information is disregarded, the argTANL identifies nearly the same number of correct predictions for the two datasets. This highlights the critical impact of label granularity on performance and underscores the necessity for a sophisticated approach capable of managing fine-grained label semantics.
@@ V-C Coarse-Grained vs Fine-Grained Label Semantics
V-D Performance Analysis Based on the Length of the Input
@@
V-E Handling Relations of Different Depth
In this section, we analyze ARs where the depth of ACs varies. Consider an arbitrary argumentative paragraph with multiple ACs, such as $t_{1}^{c}$, $t_{2}^{c}$, $t_{3}^{c}$, $t_{4}^{c}$, and $t_{5}^{c}$. These ACs can form a chain like $t_{1}^{c}\rightarrow t_{2}^{c}\rightarrow t_{3}^{c}\rightarrow t_{4}^{c}\rightarrow t_{5}^{c}$, where $t_{5}^{c}$ serves as the root. Each subsequent AC is connected to the previous one through a unique AR. This forms a chain of depth 4, as there are four distinct ARs linking these five ACs. The depth of the chain increases incrementally from the root, starting at level 1. For example, $t_{4}^{c}\rightarrow t_{5}^{c}$ is a level 1 AR, while $t_{3}^{c}\rightarrow t_{4}^{c}$ represents a level 2 AR, and so forth. In such cases, distinguishing between the ARs of closely linked ACs becomes challenging. Since all ACs in the chain are indirectly related through the root, there is a high probability of incorrectly identifying a relation, such as $t_{3}^{c}\rightarrow t_{5}^{c}$, instead of the true AR $t_{4}^{c}\rightarrow t_{5}^{c}$. Additionally, another challenge is that a single argumentative structure may contain multiple chains of different depths. To analyze this further, we use the AAE dataset, which has a tree structure with chains of maximum depth 3, making it easier to analyze the performance across chains of varying depths. Figure 8 suggests that, as the depth of the chains increases, the performance of all methods for detecting ARs decreases. Among the proposed methods, the ME-argTANL consistently outperforms the others across chains of varying depths. This clearly demonstrates the superior ability of the ME-argTANL framework to accurately identify ARs, regardless of the depth of the chains. Notably, in Figure 8, we consider both Level 1 and Level 2 ARs for the depth 2 chain analysis. Similarly, Level 1, Level 2, and Level 3 ARs are considered for the depth 3 chain analysis.
Figure 8: Percentage of correct relations for different depth chains in the ACRE task
Figure 9: Level-wise Relations (in Percentage) within Chains of Depth 2 (top) and Depth 3 (bottom) in the ACRE Task. For Chains of Depth 2, each grouped bar denotes level 1 and level 2 relations, respectively. For Chains of Depth 3, the first bar denotes level 1 relations, and the second bar represents all relations of level 2 and above.
Additionally, we analyze ARs for chains with Depths of 2 and 3 separately. For chains with a depth of 2, there are two possible levels of ARs: Level 1 and Level 2. Similarly, for chains with a depth of 3, three levels of ARs are possible: Level 1, Level 2, and Level 3. We level-wise categorize the ARs into two groups: (i) Level 1 ARs and (ii) Level 2 and above ARs. Figure 9 shows that all models effectively detect Level 1 ARs in both Depth 2 and Depth 3 chains. Among these models, ME-argTANL performs best for Depth 2 chains, while D-MKT leads for Depth 3 chains. However, for Level 2 ARs in Depth 2 chains, the performance of all models drops significantly, with the exception of D-MKT, which shows a smaller decrease compared to the other models. Similarly, for Depth 3 chains, all models exhibit reduced performance in detecting both Level 2 and Level 3 ARs. This suggests that while most models perform well in detecting ARs adjacent to the root AC, their performance significantly declines for higher-level ARs that are far from the root.
V-F Ablation with the Abbreviated argTANL framework
@@
TABLE VI: Performance comparison between argTANL and Abbreviated argTANL formulations.
@@
Experiment results in Table VI suggest that, for the AAE, AAE-FG, and CDCP datasets, the Abbreviated formulation achieved impressive average reductions in output sequence length of 23.56%, 23.01%, and 16.45%, respectively. Also, this formulation resulted in competitive performance for the ACE task, showcasing its efficiency. While the ARC task experienced a noticeable decline in performance across all datasets with this formulation, the substantial reduction in output length represents a compelling trade-off. This suggests that this formulation effectively balances performance and efficiency, capturing essential information with shorter spans in the form of “IDs”. However, for tasks that demand a deeper understanding of relational context, the argTANL framework, with its longer spans, remains superior. Overall, the Abbreviated formulation presents pragmatic and resourceful insights, offering significant benefits in terms of output length reduction while maintaining competitive performance for certain tasks.
V-G Error Analysis
TABLE VII: Error analysis for the ACRE task. IT, IC, and IF refer to Invalid Token, Invalid Component, and Invalid Format respectively.
@@ V-G1 Errors Regarding the Generated ANL
The proposed method sometimes produces invalid ANLs as the generation is not fully controllable. We identified the following three major types of ANL-related erroneous generation: (i) Invalid Token: The generated ANL consists of some out-of-vocabulary tokens or out-of-context text spans (Hallucinations). (ii) Invalid Format: The invalid ANL format includes mismatched brackets, symbols, or corrupted text. (iii) Invalid Component: The tail component connected with the relation in ANL is invalid if it is a span of text from the non-component regions. Importantly, these errors are not very frequent, and erroneous generations are discarded as negative results without undergoing any additional post-processing. Among all the datasets, D-MKT is the most effective in generating error-free ANL, whereas E-MKT produces the highest number of errors for the ACRE task.
TABLE VIII: Percentage of increase in predicted ACs and ARs of the ACRE task due to split or merged generation of ACs
Dataset Model Increased ACs Increased ARs
$T5_{argTANL}$ 13.88% 21.21%
$T5_{E-MKT}$ 13.79% 22.15%
AAE $T5_{A-MKT}$ 14.78% 23.01%
$T5_{SM-MKT}$ 13.73% 21.39%
$T5_{D-MKT}$ 13.91% 22.24%
$T5_{ME-argTANL}$ 13.57% 21.69%
$T5_{argTANL}$ 18.85% 27.54%
$T5_{E-MKT}$ 20.62% 30.29%
AAE-FG $T5_{A-MKT}$ 19.96% 29.08%
$T5_{SM-MKT}$ 19.79% 28.22%
$T5_{D-MKT}$ 19.93% 28.46%
$T5_{ME-argTANL}$ 19.60% 28.53%
$T5_{argTANL}$ 10.97% 57.71%
$T5_{E-MKT}$ 13.34% 93.20%
CDCP $T5_{A-MKT}$ 12.41% 60.26%
$T5_{SM-MKT}$ 11.69% 54.71%
$T5_{D-MKT}$ 10.58% 45.45%
V-G2 Splitted or Merged Component Boundaries
Our proposed approach encounters issues with split or merged component boundaries. If two components have boundaries defined as $(c_{i}^{s},c_{i}^{e})$ and $(c_{j}^{s},c_{j}^{e})$, they can be incorrectly merged into a single boundary $(c_{i}^{s},c_{j}^{e})$ improperly, resulting in inaccuracies. The vice-versa also happens when a single component gets split into two or more pieces, resulting in an increasing number of components compared to the number of gold components. This, in turn, propagates the error for the ARC task, increasing possible permutations for the relationships between two components. Table VIII shows the resultant increased components and relations for different frameworks for all the datasets. For both versions of the AAE dataset, argTANL and ME-argTANL demonstrate greater efficiency in handling such errors. Whereas, for the CDCP dataset, D-MKT proves superior in managing these errors.
VI Related Work
VI-A Argument Mining
Most of the prior studies have focused on only a subset of the four AM sub-tasks. However, recent works [31, 13] are focusing more on the joint formulation in an end-to-end manner. A pipelined approach for ACE and ARC one after another was followed by [27], where error propagation was optimized by performing joint inference using Integer Linear Programming (ILP). The end-to-end AM task was reformulated in four different ways by [28]: sequence tagging, dependency parsing, multi-task tagging, and relation extraction. A biaffine network-based dependency parsing for end-to-end AM was proposed by [31]. Dataset scarcity in AM literature was identified by [7], and a cross-corpora multi-task formulation with a span-biaffine architecture was proposed. A span classifier generates BIO tags of spans, and using average pooling, it generates span representations. Within the generative paradigm, the task was framed as a text-to-sequence generation task by [13]. In this generative framework, an array-like sequence is generated, consisting of AC and AR types with the start and end indices of AC spans. Recently, the TANL framework was also adopted by [35] to solve end-to-end AM as text-to-text generation. However, the extensive study of markers and the different ANL formats was unexplored in this work. Our work aims to fill this research gap.
VI-B Markers
The significance of different types of markers across various NLP tasks has been explored in the literature [21]. Several studies have shown that markers are crucial signals for ADUs [37]. Markers were used as a lexical feature for classifying argument components with a feature-based multiclass classification by [22]. Later, 1131 argumentative markers from different datasets were extracted by [20] to check the efficacy of AM tasks. An improved span representation utilizing the information of extracted markers was proposed. The contribution of markers for AM tasks in the Reddit social discussion thread was also explored by [37]. Sixty-nine Reddit-specific markers were extracted, and selective masked language modeling (sMLM) was performed by masking those markers for domain adaptation. Later, the resultant model was used for Argument Component Identification. A template-based approach was designed to predict the marker-like tokens in masked positions to predict the Relation Type between given components. However, within the current body of literature, the exploration of markers for AM tasks within a generative paradigm, employing an end-to-end AM framework, remains unexplored.
VI-C Augmented Natural Language (ANL) & Generative Paradigm
With the recent development of generative methods, most NLP tasks are being reformulated as generation problems. Generating label-augmented text (a.k.a. ANL) is one among various generative formulation strategies. It has been applied for several NLP tasks like NER [38], sentiment analysis [39], and relation extraction [40]. Recently, ANL was applied by [17] to perform various structured prediction tasks like joint entity and relation extraction, event argument extraction, and coreference resolution, by framing them as generative text-to-text translation problems.
@@ VII Conclusion
This work introduces proposes argTANL framework, a generative approach to AM. We demonstrate a text-to-text generation model could effectively handle the complexities of End-to-End AM tasks. We further enhance the argTANL framework through Marker-Based Fine-Tuning, which utilizes argumentative markers and DMs to improve the performance. This fine-tuning step enables the model to better recognize and extract ACs by learning the marker information. Additionally, we propose ME-argTANL framework, which directly integrates marker knowledge into the output sequence. ME-argTANL framework simultaneously identifies markers and argumentative structures. This dual-task approach yields superior performance compared to traditional methods. Through extensive experimentation on multiple AM benchmarks, we validate the proposed approaches argTANL, Marker-Based Fine-Tuned argTANL, and ME-argTANL by showcasing strong performance. We also explore various aspects, such as label granularity, input length, and relation depth, offering insights into the strengths and limitations of the proposed approach. The ablation study on Abbreviated argTANL provides additional evidence of the efficiency and effectiveness of our framework, especially in managing longer sequences. Thus, our research highlights the potential of generative models in the AM domain. This also paves the way for future research in integrating domain-specific knowledge into generative models for complex NLP tasks.
VIII Limitations and Future Scope
Despite the advancements and competitive performance achieved by the ME-argTANL framework, several limitations remain that warrant further exploration. The effectiveness of our model closely ties to the presence and quality of argumentative markers and DMs in the training corpus. In domains where such markers are absent or explicit, further investigation is required to improve the model’s performance in such scenarios. We conduct the experiments in a single-corpus setting. Investigating the performance of our generative method in a multi-corpus setup can provide valuable insights. Additionally, exploring more advanced fine-tuning strategies beyond the four two-step strategies can yield higher performance gains. Our approach uses the standard AM benchmarks for fine-tuning marker-based strategies. Evaluating the system’s robustness in noisy real-world scenarios would also be a compelling avenue for future research. While our experiments primarily utilize the T5 model, exploring other large language models such as BART, Llama, and Mistral could offer additional perspectives on their performance for this task.

### v3

@@ IV-C3 Marker Knowledge Transfer through Encoding (E-MKT)
IV-C4 Discourse Marker Knowledge Transfer (D-MKT)
@@ IV-D The ME-argTANL Framework
V Experimental Setup
V-A Datasets
We consider three standard datasets to evaluate the proposed methods as follows:
Argument Annotated Essay (AAE)[25]: This dataset contains 402 student essays annotated at the segment (span) level. Every essay is divided into multiple paragraphs. A total of 1833 paragraphs are annotated with three AC types, $T^{c}=\{\text{{Claim, MajorClaim, Premise}}\}$, and two AR types, $T^{r}=\{\text{{Support, Attack}}\}$. AAE contains a large number of argumentative markers, with almost every AC beginning with one. We extract argumentative markers only from this dataset and transfer this knowledge to all experiments.
Fine-Grained Argument Annotated Essay (AAE-FG)[29]: Recently, the AAE dataset is annotated in a more fine-grained manner by [29]. The Major Claim and Claim classes are divided into Fact, Value, Policy. And the Premise class is divided into Common Ground, Testimony, Hypothetical Instance, Statistics, Real Example and Others. So, it contains a total of nine AC types and the ARs remain unchanged.
Consumer Debt Collection Practices (CDCP)[2]: This dataset contains 731 user comments collected from the Consumer Financial Protection Bureau (CFPB) website. It contains five AC types, $T^{c}=\{\text{{Fact, Testimony, Reference, Policy, Value}}\}$, and two AR types, $T^{r}=\{\text{{Reason, Evidence}}\}$.
@@ V-A Datasets
TABLE II: Experiment results on ACRE task with comparable baselines. Best scores are marked in bold. * indicates the baseline results we produced by running their open-source code.
@@
V-B Training Details
We conducted all experiments using the T5-Base model [31] on Nvidia A100 GPUs, with each reported result averaged over 5 runs [^1] The implementation will be made available upon acceptance.. The hyperparameter settings for the argTANL framework are as follows: For the CDCP dataset, a batch size of 4 is used, while for AAE and AAE-FG, the batch size is set to 8. Across all datasets, we utilize the AdamW optimizer with a learning rate of 0.0005. The maximum input length is set to 1024 tokens for CDCP and 512 tokens for AAE and AAE-FG. In the end-to-end and component-only training scenarios, we run for more than 10000 steps, with checkpoints saved every 200 steps. During inference, beam search is used with a beam length of 8 across all datasets. For all strategies (A-MKT, SM-MKT, D-MKT, and E-MKT) of Marker-Based Fine-Tuning, we use a batch size of 16, the AdamW optimizer with a learning rate of 0.0005, a maximum input length of 512 tokens, and 10 epochs of training. We use the following libraries: (i) TANL framework [^2] https://github.com/amazon-science/tanl, and (ii) HuggingFace’s Transformers [^3] https://github.com/huggingface.
V-C Baseline Methods
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
  7. 7.
  8. 8.
  9. 9.
  10. 10.
  11. 11.
    DENIM [37]: A discourse structure aware generative framework with multitask prompt-tuning.
V-D Performance Metrics for Evaluation
After generating the target ANL, we post-process the sequence by removing the symbol tokens to get a cleaned text. Finally, for a comprehensive evaluation, AC and AR tuples are created, including their types and corresponding boundaries. Following [22], we evaluate the results with micro-F1 score for both ACE and ACRE tasks, where an exact match with the gold label is considered as a true label.
VI Results and Discussion
@@
Table II compares the performance of the proposed frameworks with the baselines on the ACRE task in terms of micro F1-scores for component classification and relation classification. For the AAE dataset, $T5_{ME-argTANL}$ turns out to be the best among the proposed frameworks, surpassing argTANL and all the MFS-based strategies, for component classification. While $T5_{ME-argTANL}$ outperforms several baselines for component and relation classifications, it gave a competitive performance to the current state-of-the-art models for component classification. We observe a similar performance on the AAE-FG dataset. For the AAE-FG dataset, $T5_{ME-argTANL}$ surpasses the current state-of-the-art result for component classification.
@@
TABLE III: Performance difference on the ACE task: Component-Only vs. End-to-End variants. Among all techniques, minimum differences between these two variants are marked in bold.
@@
VI-A Component Only vs End-to-End Variants
Table III compares the performance of models on the ACE task in the two formulations. Analyzing the argumentative structure of AAE and AAE-FG datasets reveals that argument components always form a tree, indicating a single parent for each child. In this well-structured dataset, the performance of the models drops in the Component-only variant compared to the end-to-end variant. The CDCP dataset contains many isolated components. It does not necessarily form a tree but can form a graph (e.g., $t_{1}^{c}$$t_{2}^{c}$, $t_{2}^{c}$$t_{3}^{c}$, and a transitive relation $t_{1}^{c}$$t_{3}^{c}$). We observe the similar performance of models, except E-MKT, in the two variants for CDCP corpus. This result indicates that ACs and ARs benefit from mutual feature information, suggesting synergy in an end-to-end setup. For Component-only variant, the D-MKT strategy proves to be beneficial for all the datasets. When comparing the compatibility of the E-MKT with other MFS strategies, it appears more suitable for the Component-only variant rather than the end-to-end for the ACE task. It exhibits the minimum performance difference in the Component-only setup for both versions of the AAE corpus and even shows performance improvements for the CDCP corpus.
VI-B Performance Analysis Based on the Length of the Input
@@
VI-C Ablation with the Abbreviated argTANL framework
@@
TABLE IV: Performance comparison between argTANL and Abbreviated argTANL formulations.
@@
Experiment results in Table IV suggest that, for the AAE, AAE-FG, and CDCP datasets, the Abbreviated formulation achieved impressive average reductions in output sequence length of 23.56%, 23.01%, and 16.45%, respectively. Also, this formulation resulted in competitive performance for the ACE task, showcasing its efficiency. While the ARC task experienced a noticeable decline in performance across all datasets with this formulation, the substantial reduction in output length represents a compelling trade-off. This suggests that this formulation effectively balances performance and efficiency, capturing essential information with shorter spans in the form of “IDs”. However, for tasks that demand a deeper understanding of relational context, the argTANL framework, with its longer spans, remains superior. Overall, the Abbreviated formulation presents pragmatic and resourceful insights, offering significant benefits in terms of output length reduction while maintaining competitive performance for certain tasks.
VI-D Error Analysis
TABLE V: Error analysis for the ACRE task. IT, IC, and IF refer to Invalid Token, Invalid Component, and Invalid Format respectively.
@@ VI-D Error Analysis
The proposed method sometimes produces invalid ANLs as the generation is not fully controllable. We identified the following three major types of ANL-related erroneous generation (see Table V): (i) Invalid Token: The generated ANL consists of some out-of-vocabulary tokens or out-of-context text spans (Hallucinations). (ii) Invalid Format: The invalid ANL format includes mismatched brackets, symbols, or corrupted text. (iii) Invalid Component: The tail component connected with the relation in ANL is invalid if it is a span of text from the non-component regions. Importantly, these errors are not very frequent, and erroneous generations are discarded as negative results without undergoing any additional post-processing. Among all the datasets, D-MKT is the most effective in generating error-free ANL, whereas E-MKT produces the highest number of errors for the ACRE task.
@@ VII Conclusion
This work introduces proposes argTANL framework, a generative approach to AM. We demonstrate a text-to-text generation model could effectively handle the complexities of End-to-End AM tasks. We further enhance the argTANL framework through Marker-Based Fine-Tuning, which utilizes argumentative markers and DMs to improve the performance. This fine-tuning step enables the model to better recognize and extract ACs by learning the marker information. Additionally, we propose ME-argTANL framework, which directly integrates marker knowledge into the output sequence. ME-argTANL framework simultaneously identifies markers and argumentative structures. This dual-task approach yields superior performance compared to traditional methods. Through extensive experimentation on multiple AM benchmarks, we validate the proposed approaches argTANL, Marker-Based Fine-Tuned argTANL, and ME-argTANL by showcasing strong performance. The ablation study on Abbreviated argTANL provides additional evidence of the efficiency and effectiveness of our framework, especially in managing longer sequences. Thus, our research highlights the potential of generative models in the AM domain. This also paves the way for future research in integrating domain-specific knowledge into generative models for complex NLP tasks.